### v1

# Cross-Service Threat Intelligence in LLM Services using Privacy-Preserving Fingerprints

Abstract
The widespread deployment of LLMs across enterprise services has created a critical security blind spot. Organizations operate multiple LLM services handling billions of queries daily, yet regulatory compliance boundaries prevent these services from sharing threat intelligence about prompt injection attacks, the top security risk for LLMs. When an attack is detected in one service, the same threat may persist undetected in others for months, as privacy regulations prohibit sharing user prompts across compliance boundaries.
We present BinaryShield, the first privacy-preserving threat intelligence system that enables secure sharing of attack fingerprints across compliance boundaries. BinaryShield transforms suspicious prompts through a unique pipeline combining PII redaction, semantic embedding, binary quantization, and randomized response mechanism to potentially generate non-invertible fingerprints that preserve attack patterns while providing privacy. Our evaluations demonstrate that BinaryShield achieves an F1-score of 0.94, significantly outperforming SimHash (0.77), the privacy-preserving baseline, while achieving 64x storage reduction and 38x faster similarity search compared to dense embeddings.
1 Introduction
The rapid adoption of Large Language Models (LLMs) in digital services is driving a paradigm shift in human-computer interaction. Companies now operate many LLM-based services across diverse domains, collectively handling billions of queries each day. Typically, organizations like Microsoft maintain multiple, logically isolated (siloed) LLM services, such as enterprise AI assistants, consumer chat applications, API based LLM services, and developer-focused coding agents. These services are siloed because they are tailored to meet distinct business objectives and end-user requirements, necessitating separate operational boundaries. Each service has distinct model stacks, logging pipelines, and compliance boundaries with strong internal data governance policies to protect user data and privacy. In addition to internal privacy regulations, these services are also subject to governmental regulations such as GDPR and HIPAA, which mandate strict data handling practices to protect user privacy and sensitive information. While this siloed architecture is crucial for maintaining operational independence, it has an unintended consequence: fragmented security telemetry and a weakened collective defense. With little correlation or data sharing between services, organizations are left with a disjointed security posture as each service operates independently, responding to threats in isolation. This fragmentation creates a significant challenge for incident response teams, as they lack a unified view of the threat landscape across their LLM services.
[…]
Motivation and Problem Statement. The current defense paradigm is compartmentalized, with each service responding independently to detected attacks. When a new attack vector is identified (typically through media reports, user complaints, or post-breach analysis) security teams patch their systems (e.g., update prompts classifiers). However, this knowledge remains confined to the affected service. Furthermore, attacks are inevitable because existing defenses are probabilistic and can be bypassed (Debenedetti et al., 2025; Costa et al., 2025). Thus, organizations need retrospective systems to pinpoint the perpetrator and timing of incidents.
[…]
Our Contributions. We propose the concept of threat intelligence across compliance boundaries in LLM services. We materialize this idea in BinaryShield, the first privacy-preserving fingerprinting technique for prompt injection attacks that enables secure information sharing between otherwise siloed services. When a potential threat (prompt injection) is detected within any service by its defense mechanism, BinaryShield generates a non-reversible, lightweight fingerprint that distills the semantic essence of the malicious prompt, making it possible to share actionable threat signals while preserving privacy.
Our approach balances the need for threat intelligence sharing with user privacy and regulatory compliance. BinaryShield achieves this through a multi-stage process that progressively removes identifying information while retaining the essential semantic characteristics for effective threat detection. The process begins with personally identifiable information (PII) redaction, where BinaryShield removes sensitive data such as social security numbers, names, and other identifying markers from input prompts. This initial privacy layer is crucial for protecting users whose prompts may be incorrectly flagged as malicious by automated defenses. However, PII redaction alone is insufficient, as the remaining text may still contain contextual information that could compromise user privacy when shared externally. To capture the semantic essence of potentially malicious prompts, BinaryShield generates high-dimensional embeddings from the redacted text. These embeddings, which represent text as dense floating-point vectors, achieve state-of-the-art performance in natural language processing tasks by encoding semantic relationships in vector space. However, embeddings are not secure for sharing across compliance boundaries. Recent research demonstrates that embeddings pose significant privacy risks, as they can allow adversaries to reconstruct original input text (Walsh, ; Li et al., 2023; Tragoudaras et al., 2025; OWASP2025, 2025). Moreover, traditional embeddings present practical challenges for large-scale threat intelligence system. Each dimension typically requires 32 bits of storage, and similarity computations using cosine similarity or dot products demand substantial computational resources, often necessitating GPU acceleration for efficient threat matching.
BinaryShield addresses these fundamental challenges through two innovative non-reversible transformations by exploiting insights from quantization and differential privacy. First, BinaryShield performs binary quantization on the floating-point embeddings, converting each dimension to a single bit based on its sign, assigning 1 for positive values and 0 for negative values. This quantization achieves remarkable efficiency gains, reducing storage requirements while simultaneously enhancing privacy. The key insight underlying this approach is that by discarding magnitude information and retaining only directional information, we make prompt reconstruction exponentially more difficult while preserving sufficient semantic structure for threat sharing and detection. To render original prompt reconstruction practically impossible, BinaryShield applies differential privacy to the binary embeddings. Leveraging the insights from the principle of randomized response (Warner, 1965; Erlingsson et al., 2014; Wang et al., 2017), BinaryShield independently flips each bit of the binary vector with a precisely calibrated probability. This noise addition fundamentally alters the binary vector representation while preserving its calibrated utility for threat correlation, ensuring that adversaries cannot reverse-engineer the original prompt from the shared fingerprint. The resulting binary vector is efficient in storage and computation, enabling rapid Hamming distance similarity checks. More importantly, after these privacy-preserving transformations, the resulting binary vector becomes safe to share across compliance boundaries, enabling effective cross-service threat intelligence without compromising user privacy or regulatory compliance.
Figure 1: BinaryShield system design. Suspicious prompts are processed within the compliance boundary to generate privacy-preserving fingerprints, which are then shared across services for collaborative threat detection.
Evaluations. We conduct comprehensive evaluation of BinaryShield spanning threat correlation effectiveness, privacy calibration, scalability, and operational efficiency (Section 3). BinaryShield significantly outperforms privacy-preserving baseline across attack variants (17-point F1 advantage on sophisticated attacks) and exhibits smooth privacy-utility trade-offs with theoretical noise alignment. In real-world enterprise-scale scenarios, BinaryShield achieves 79.2% threat correlation accuracy (93% of non-private baseline performance) alongside efficiency gains (38x search speedup, 64x storage reduction).
2 BinaryShield Design
The proliferation of LLM-based services within organizations creates a fundamental tension between security effectiveness and privacy compliance. While traditional malware defense systems successfully share threat signatures across boundaries, no comparable mechanism exists for natural language-based attacks on LLMs. BinaryShield addresses this gap by enabling organizations to share actionable threat intelligence about prompt injection attacks without violating privacy regulations or exposing sensitive user data.
The system is designed to achieve four primary objectives: (1) Semantic Preservation: capture the essential characteristics of malicious prompts that enable cross-service threat correlation; (2) Privacy Guarantee: ensure computational non-invertibility such that original prompts cannot be reconstructed from shared fingerprints hence upholding compliance with internal and government regulations (i.e. GDPR); (3) Computational Efficiency: enable efficient fingerprint generation and search; and (4) Operational Compliance: maintain strict separation of user data across compliance boundaries while enabling collaborative defense. To this end, BinaryShield transforms potentially malicious prompts through a series of privacy-preserving operations, ultimately producing compact binary fingerprints that can be safely shared between services while maintaining regulatory compliance. The key insight underlying our approach is that prompt injection attacks, despite surface-level variations, share semantic patterns that can be captured and compared without revealing the original content. By combining techniques from data anonymization, natural language processing, quantization and differential privacy, we create non-reversible fingerprints that preserve these essential semantic characteristics while providing privacy guarantees against reconstruction attacks.
[…]
Figure 2: BinaryShield approach for privacy-preserving fingerprint generation for cross-service threat intelligence sharing. The pipeline progressively transforms potential prompt injections through PII redaction, semantic embedding, binary quantization, and differential privacy to produce shareable fingerprints that preserve privacy while enabling threat correlation across compliance boundaries.
High-Level Architecture.
Figure 1 illustrates the complete BinaryShield system architecture. An attacker issues semantically related prompts ($x,y,z$) to the services of an enterprise organization that reside in separate compliance boundaries (Figure 1-➊). Each service applies their own defense mechanism ($D1$, $D2$, and $D3$) to incoming prompts. Service-2 flags a prompt-injection attempt (Figure 1-➋). After this step, BinaryShield’s fingerprinting pipeline is invoked in the service’s compliance boundary (Figure 1-➌). At this stage, the attack log is separated into two components. The first and the most critical component is private information: this contains the sensitive content of the prompt itself, holding the essential details of the attack. Although critical for accurate event correlation, this data must remain within the service’s compliance boundary to safeguard privacy due to regulatory requirements. The suspicious prompt then undergoes a series of privacy-preserving transformations to generate a non-reversible privacy fingerprint (hash) within the attacked service’s compliance boundary (Figure 1-➌). A detailed description of the privacy fingerprint generation process appears in Section 2.1. The second part consists of non-private system metadata, such as the tools used during response generation and the geographical region. Since this metadata contains no sensitive data, it can be safely shared among peer services as part of complete signature. Sharing these features enhances attack correlation by providing complementary evidence to surface relevant patterns pointing to attack source and channel, which helps reduce false positives and prioritize alerts. For instance, matching system metadata across services, such as both accessing an email tool, significantly increases the confidence in the correlation of potential attacks. Concatenating both parts forms a composite attack fingerprint that discloses no readable text (Figure 1-➍) This encoded hash is then securely broadcast to peer services for threat correlation (Figure 1-➎). Note that peer services (i.e., Services 1 and 3) use the exact same technique to fingerprint their local private queries and system metadata as used in Step ➌ of Figure 1. For instance, after searching their logs, Service 1 and Service 3 independently find two and one similar attacks, respectively (Figure 1-➏). Policy-driven response actions are then executed for each service based on the correlation results (Figure 1-➐). Note that there is currently no public dataset available for system metadata. Therefore, our evaluation of BinaryShield focuses exclusively on private information (prompt) fingerprinting. Nonetheless, including system metadata features in the overall design demonstrates the full architectural intent of BinaryShield and highlights additional layers of security and context.
2.1 BinaryShield One-Way Fingerprint Generation Design
1
Input: Suspicious prompt $y$; system metadata $m$; privacy parameter $\alpha$; set of peer services $\mathcal{S}$
Output: Privacy-preserving fingerprint $f$ broadcast to $\mathcal{S}$; correlated alerts at peers
2
$y_{redacted}\leftarrow$ PII_Redact($y$) ;
  // Remove all PII, replace with placeholders
3
$\mathbf{e}\leftarrow$ Embed($y_{redacted}$) ;
  // Compute semantic embedding vector
4
$\mathbf{b}\leftarrow[]$ ;
  // Initialize empty binary vector
5 for $i\leftarrow 1$ to $d$ do
6 $b_{i}\leftarrow\begin{cases}1&\text{if }e_{i}>0\\ 0&\text{otherwise}\end{cases}$ ;
7   Append $b_{i}$ to $\mathbf{b}$ ;
8
$p_{\text{keep}}\leftarrow\frac{e^{\alpha}}{e^{\alpha}+1}$ ;
  // Probability of keeping the true bit
$\tilde{\mathbf{b}}\leftarrow$[] ;
  // Initialize empty privatized vector
9 for $i\leftarrow 1$ to $d$ do
10   Sample $u_{i}\sim\mathrm{Uniform}(0,1)$ ;
11 if $u_{i}<p_{\text{keep}}$ then
    $\tilde{b}_{i}\leftarrow b_{i}$ ;
       // Keep the true bit
12
13 else
    $\tilde{b}_{i}\leftarrow 1-b_{i}$ ;
       // Flip the bit
14
15  Append $\tilde{b}_{i}$ to $\tilde{\mathbf{b}}$ ;
16
[…]
  // Combine privatized bits with non-private metadata
17
Broadcast($f$, $\mathcal{S}$) ;
  // Send fingerprint to all peers
18
Algorithm 1 BinaryShield: Privacy-Preserving Threat Intelligence Pipeline
We consider an adversary with several capabilities: possesses white-box access, meaning they have complete knowledge of the BinaryShield algorithm, including the embedding models, quantization methods, and privacy parameters. The adversary’s primary objective is to reconstruct original prompts from shared fingerprint, potentially exposing sensitive user information. Our system must, therefore, prevent such reconstruction while preserving utility for effective threat detection. Based on our threat model and regulatory constraints, BinaryShield must satisfy several critical requirements. First, it must ensure privacy preservation by making fingerprints computationally non-invertible, thereby preventing the reconstruction of original prompts even in the case of white-box access. Second, semantic preservation remains essential; despite privacy transformations, the fingerprints must retain enough semantic information to enable the identification of similar attack patterns. Third, the system must maintain computational efficiency, capable of processing millions of queries each day with minimal latency impact. Fourth, it must offer cross-boundary compatibility, allowing fingerprints to be shared across services with varying compliance requirements without violating relevant regulations. Finally, the system must support approximate matching, enabling similarity searches that can detect related attacks despite minor variations in the input. Figure 2 and Algorithm 1 depicts the BinaryShield pipeline for generating privacy-preserving fingerprints (hashes) of prompt injection attacks. Below, we detail each component of the pipeline, explaining how they collectively satisfy our design requirements.
2.1.1 PII Redaction
The initial stage of the BinaryShield pipeline directly targets privacy risks by identifying and removing personally identifiable information (PII) that may be present in flagged prompts. While automated defense mechanisms aim to identify malicious content, they may inadvertently flag benign prompts containing sensitive user data.
Figure 3: Distribution of PII entities in prompts flagged by BinaryShield. The most common entity include person names.
A key challenge in PII redaction is removing sensitive information while preserving the semantic intent and structure of potential attacks. Our approach addresses this by detecting and redacting structured data such as social security numbers, credit card numbers, phone numbers, email addresses, as well as person names, organizations, locations, and other identifying entities. Identified PII is replaced with generic placeholders that maintain command structure, entity relationships, and linguistic patterns indicative of injection attempts. For example, "Transfer \$5000 from John Smith’s account 123456789" becomes "Transfer [AMOUNT] from [PERSON]’s account [ACCOUNT]". This balance ensures the redacted text retains sufficient information for downstream processing while protecting user privacy from the earliest stage of the pipeline. By design, redaction is intended to be one-way and non-reversible: no original PII should survive beyond this step, ensuring compliance with privacy regulations before any downstream processing. While the effectiveness of redaction depends on the accuracy of underlying detection methods, our approach minimizes the risk of residual PII to the greatest extent possible. BinaryShield uses Microsoft’s Presidio open-source library (Microsoft, ) for PII detection and redaction. Figure 3 reports the PII entities and their respective counts that BinaryShield flags in the prompt injection dataset from (Shen et al., 2024), used in BinaryShield’s evaluation (Section 3). However, even after PII redaction, the remaining text may still contain sensitive or revealing information that cannot be safely shared across compliance boundaries. To address this, BinaryShield’s pipeline generates privacy-preserving semantic fingerprints, enabling robust threat detection without exposing any underlying content.
2.1.2 Semantic Feature Extraction
Following PII redaction, BinaryShield must capture the semantic essence of the potentially malicious prompt in a format suitable for comparison and analysis. To this end, BinaryShield employs state-of-the-art transformer-based language models to generate high-dimensional embeddings that encode the semantic relationships within the text. This transformation serves as the foundation for our semantic preservation requirement. This capture the meaning and intent of the potentially malicious prompt. The embedding process transforms the redacted text into a dense vector $\mathbf{e}\in\mathbb{R}^{d}$, where $d$ is the embedding dimension. The embedding space exhibits several properties that are crucial for threat detection. First, similar attack patterns tend to cluster together in the embedding space, regardless of surface-level variations. Second, the embeddings are robust to paraphrasing, so minor rewording or synonym substitution results in nearby embeddings, enabling the detection of attack variants.
Binary Quantization.
While semantic embeddings effectively capture prompt meaning, they pose significant challenges for privacy-preserving threat intelligence. Dense embeddings can leak information about the original text (Walsh, ; Li et al., 2023; Tragoudaras et al., 2025; OWASP2025, 2025). They also require substantial computational resources for large-scale similarity searches. Our binary quantization addresses both concerns. We transform the continuous embedding $\mathbf{e}\in\mathbb{R}^{d}$ into a binary vector $\mathbf{b}\in\{0,1\}^{d}$ using sign-based quantization:
$b_{i}=\begin{cases}1&\text{if }e_{i}>0\\ 0&\text{otherwise}\end{cases}$
where $e_{i}$ is the $i$-th dimension of the embedding vector and $b_{i}$ is the corresponding bit in the binary vector. This transformation discards magnitude information while preserving the directional (sign) information of each embedding dimension. Binary quantization provides several critical benefits: Irreversibility: The transformation from $\mathbb{R}$ to $\{0,1\}$ is many-to-one, making exact reconstruction impossible. Each bit only indicates whether the original value was positive or negative, discarding all magnitude information. Dimensionality preservation: Despite the dramatic reduction in information per dimension, the high dimensionality ($d$) of the binary embedding preserves the essential geometric properties needed for similarity detection. The sign of each embedding dimension collectively maintains sufficient information to distinguish between semantically similar and dissimilar prompts. Storage efficiency: Each dimension requires only 1 bit instead of 32 bits (float32), achieving a 32x reduction in storage requirements and even more in case of float64. Our empirical evaluation shows that binary quantization preserves semantic relationships and achieves comparable threat detection performance, while providing substantial privacy benefits (Section 3).
Figure 4: Impact of BinaryShield’s privacy parameter ($\alpha$) on binary vector quality. As $\alpha$ increases, privatized vector (red triangles) move closer to original vector (blue circles), showing improved utility while maintaining differential privacy through controlled bit-flipping (i.e., 1p1-p) noise.
2.1.3 BinaryShield’s Randomized-Response Layer
While data anonymization (PII redaction) and binary quantization significantly enhances privacy, adversaries with auxiliary information might still attempt reconstruction attacks. To close this gap, the pipeline enforces local differential privacy (LDP) on every fingerprint by applying classical randomized response (Warner, 1965; Erlingsson et al., 2014; Wang et al., 2017) independently to each bit before the fingerprint crosses the compliance boundary. Local differential privacy is particularly attractive in cross-boundary threat-intelligence settings because the guarantee is enforced before data leave the originator’s compliance perimeter; no trusted aggregator or secure enclave is required. Even an adversary with white-box knowledge of the entire pipeline and unlimited access to all shared fingerprints cannot violate the worst-case privacy bound. Moreover, in contrast to the global-DP setting, the magnitude of noise required for LDP is independent of the dataset size, allowing BinaryShield to scale to millions of daily fingerprints without ever decreasing privacy protection. Consider a $d$-bit fingerprint $\mathbf{b}=(b_{1},\dots,b_{d})\!\in\!\{0,1\}^{d}$ and let $\alpha\!>\!0$ denote the per-bit privacy budget. Classical randomized response publishes a perturbed vector $\tilde{\mathbf{b}}$ obtained by flipping each bit independently with probability 1p1-p, where
$p=\frac{e^{\alpha}}{e^{\alpha}+1}$
The mechanism releases the randomized bit.
$\tilde{b}_{i}\;=\;\begin{cases}b_{i}&\text{with probability }p,\\[6.0pt] 1-b_{i}&\text{with probability }1-p.\end{cases}$
Randomized response inevitably perturbs the underlying signal. The expected self-distance introduced by the mechanism is
$\mathbb{E}\bigl{[}\mathrm{H}(\mathbf{b},\mathbf{\tilde{b}})\bigr{]}\;=\;(1-p)d,$
where $\mathrm{H}$ denotes Hamming distance and $d$ is the fingerprint dimensionality. The Hamming distance between two binary vectors $\mathbf{b},\mathbf{\tilde{b}}\in\{0,1\}^{d}$ is defined as the number of positions at which the corresponding bits differ:
$\mathrm{H}(\mathbf{b},\mathbf{\tilde{b}})=\sum_{i=1}^{d}\mathbb{I}[b_{i}\neq\tilde{b_{i}}$]
where $\mathbb{I}[\cdot]$ is the indicator function, which equals 1 if its argument is true and 0 otherwise. Because $p$ remains close to one even for moderate $\alpha$, the added noise is sparse and, as we show experimentally in Section 3, preserves enough structure for accurate approximate matching while providing the formal privacy guarantees demanded by cross-service threat intelligence.
The randomization budget $\alpha$ serves as a control parameter governing the privacy-utility tradeoff in BinaryShield. When $\alpha$ approaches zero, the mechanism maximizes privacy protection (i.e., infinite privacy) but renders fingerprints unusable for correlation. As $\alpha$ increases, utility improves while maintaining privacy guarantees in a controlled manner. Figure 4 illustrates this relationship through PCA visualization of binary vectors before and after applying the Randomized Response mechanism (Warner, 1965). At $\alpha=0.25$, privatized vectors (red triangles) appear completely random. As $\alpha$ increases (reducing noise), these vectors progressively align closer to their original non-private counterparts (blue circles). Notably, even at $\alpha=3$, a non-zero average displacement between original and privatized vectors persists, making reconstruction of the original prompt challenging. This visualization confirms that larger $\alpha$ values preserve more of the vector’s structure while maintaining formal differential privacy guarantees through controlled bit-flipping. In practice, higher privacy budgets may be appropriate for intra-service sharing where additional security measures already exist within organizational boundaries.
2.1.4 BinaryShield’s Cross-Service Threat Correlation
When a service detects a potential prompt injection, it generates a fingerprint using the BinaryShield pipeline. For clarity, we denote the locally randomized-response fingerprint $\tilde{\mathbf{b}}\in\{0,1\}^{d}$ as $f$ when broadcasting and correlating across services (i.e., $f\approx\tilde{\mathbf{b}}$). This fingerprint is broadcast asynchronously to participating services through secure channels. Upon receiving a fingerprint, services scan their historical logs for similar instances. If matches are found, alerts are triggered and defenses can be updated proactively. To maintain privacy and compliance, only aggregate match statistics are shared, and no specific prompt content is revealed. This workflow enables rapid propagation of threat intelligence while respecting organizational boundaries. Formally, let $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{N}\}$ denote the set of $N$ services, each with its own compliance boundary. Each service $S_{i}$ maintains a log of previously observed fingerprints $\mathcal{F}_{i}=\{f_{i}^{(1)},f_{i}^{(2)},\ldots,f_{i}^{(M_{i})}\}$, where $f_{i}^{(m)}\in\{0,1\}^{d}$ is a $d$-dimensional binary fingerprint. Suppose service $S_{q}$ detects a suspicious prompt and generates a fingerprint $f_{q}\in\{0,1\}^{d}$ using the BinaryShield pipeline. This fingerprint is securely broadcast to all other services $S_{j}$ ($j\neq q$). Each recipient service $S_{j}$ performs the following search:
$\mathcal{M}_{j}(f_{q})=\left\{f_{j}^{(m)}\in\mathcal{F}_{j}\;\middle|\;\mathrm{H}(f_{q},f_{j}^{(m)})\leq\tau\right\}$
where $\mathcal{M}_{j}(f_{q})$ is the set of fingerprints belonging to service $S_{j}$ (i.e., fingerprints in $\mathcal{F}_{j}$) that match $f_{q}$ within a Hamming distance threshold $\tau$, and $\mathrm{H}(\cdot,\cdot)$ is the Hamming distance as previously defined. This protocol enables rapid, privacy-preserving propagation of threat intelligence across all $N$ services. This protocol generalizes to any number of services. For example, if Service 2 ($S_{2}$) generates a fingerprint $f_{2}$, it broadcasts $f_{2}$ to Services 1 and 3 ($S_{1}$, $S_{3}$), which independently search their logs for similar fingerprints. If $S_{1}$ finds two matches and $S_{3}$ finds one, these services take action based on their policy and only the counts are reported back, enabling rapid, privacy-preserving propagation of threat intelligence across organizational boundaries.
Overall, when a match is found below a calibrated threshold, the service can take protective actions such as updating classifier parameters, flagging suspicious activity, or blocking potential attacks. Services can subscribe to threat intelligence feeds from peer services, automatically incorporating new attack patterns into their defense mechanisms.
Summary.
BinaryShield offers a scalable and privacy-preserving solution for sharing prompt injection threat intelligence across organizational and compliance boundaries. By combining PII redaction, semantic embedding, binary quantization, and differential privacy, the system achieves strong privacy protections without sacrificing detection performance or scalability. The modular architecture allows for privacy parameters to be tuned to specific requirements, while efficient processing supports production-scale deployments. As large language model adoption expands, BinaryShield establishes a foundation for collaborative and adaptive defenses against evolving prompt injection attacks.
3 Evaluation
We conduct a comprehensive evaluation of BinaryShield to assess its effectiveness in privacy-preserving threat intelligence sharing for prompt injection attacks. Our evaluation addresses key research questions across diverse, systematically generated attack variants. BinaryShield’s evaluation is structured around the following research questions:
  • RQ1 (Detection Effectiveness). How accurately does BinaryShield detect prompt injection attacks across systematically generated adversarial variants, from simple word substitutions to comprehensive semantic paraphrasing, compared to the SimHash baseline?
  • RQ2 (Privacy-Utility Trade-off). What is the quantitative relationship between BinaryShield’s randomized response mechanism ($\alpha$) and threat detection utility?
  • RQ3 (Noise Calibration & Predictability). Does the empirical self-Hamming distortion introduced by randomized response match the theoretical (1p)d(1-p)d curve across $\alpha$, letting us confidently choose a privacy setting?
  • RQ4 (Scalability). Does BinaryShield maintain consistent detection accuracy as corpus size scales from thousands to hundreds of thousands of entries, reflecting realistic enterprise deployment scenarios where malicious prompts are sparse within benign prompts?
  • RQ5 (Computational Efficiency). What are the computational and storage efficiency gains of BinaryShield’s binary fingerprints compared to dense embeddings, and do these improvements enable practical real-time threat correlation at enterprise scale?
3.1 Evaluation Settings
3.1.1 Datasets
We evaluate BinaryShield using the recent prompt injection dataset from (Shen et al., 2024), which contains real-world attack and benign prompts. From this, we construct a comprehensive evaluation set simulating adversarial scenarios, enabling controlled and realistic assessment of threat intelligence in LLM services.
Given the inherent challenge of obtaining large-scale labeled datasets of semantically similar prompt injection variants, we develop a systematic methodology for generating controlled attack modifications to validate threat intelligence. Our synthetic data generation process creates two primary types of variants: Word-Flipping Variants: For each adversarial prompt, we randomly select $x$ words (where $x\in\{1,3,5,10,20\}$) longer than four characters and replace them with semantically equivalent synonyms using GPT-4o. This simulates subtle adversarial modifications that maintain attack intent while altering surface-level tokens. The language model is guided using the structured prompt: "Understand this adversarial prompt: <original prompt>. Now generate synonyms for the following words: <words to change>. Return the JSON with the field ’words_changes_to’ containing the synonyms in order." Paraphrase Variants: We generate comprehensive paraphrased versions of prompts using GPT-4o while preserving the underlying attack intent and malicious purpose. The generation prompt instructs the model: "Paraphrase the following adversarial prompt while maintaining the attack intent and purpose. Original prompt: <original prompt>. Return the JSON with the field ’generated_response’." Benign Pairs: Benign prompts are systematically paired such that each pair consists of two semantically unrelated prompts, ensuring they serve as appropriate negative examples for similarity detection algorithms (e.g., SimHash and BinaryShield). The resulting dataset undergoes rigorous filtering to ensure quality and consistency (e.g., remove duplicates, empty prompts). The final dataset is balanced with equal numbers of attack and benign pairs, with attack pairs labeled as 1 and benign pairs labeled as 0. This balanced approach ensures unbiased evaluation across both positive and negative cases. Our synthetic generation methodology enables systematic evaluation across varying levels of attack sophistication, from minimal single-word modifications to comprehensive paraphrasing that significantly alters surface structure while preserving malicious intent. This controlled approach allows us to assess algorithm robustness across the spectrum of adversarial modifications likely encountered in real-world deployment scenarios.
3.1.2 Baseline
Our evaluation compares two primary approaches for prompt similarity detection BinaryShield and SimHash. BinaryShield utilizes ModernBert  for semantic feature extraction, followed by sign-based binary quantization and randomized response for local differential privacy. We also evaluate it with OpenAI embedding model to demonstrate the flexibility of BinaryShield in supporting both open source and closed source models. SimHash is a widely-adopted locality-sensitive hashing algorithm that serves as our primary baseline. SimHash generates compact binary fingerprints and has been extensively used for near-duplicate detection in large-scale text corpora (Manku et al., 2007).
3.1.3 Evaluation Metrics
We employ a comprehensive set of metrics to assess both detection performance and privacy-utility tradeoffs. For detection performance, we use precision, recall, accuracy, and F1-score across varying similarity thresholds to evaluate attack detection accuracy. Threshold analysis is performed using precision-recall curves to identify the optimal threshold for Hamming distance and to understand how performance is affected by threshold selection, as discussed in (Manku et al., 2007). We also include confusion matrices for detailed analysis of true positives (TP), false positives (FP), true negatives (TN), and false negatives (FN) at optimal thresholds.
(a) Precision-recall curves across attack variants.
(b) Confusion matrices at optimal thresholds.
Figure 5: BinaryShield performance analysis across attack variants with $\alpha=2.0$. (a) Precision-recall curves demonstrate consistent performance across variants. (b) Confusion matrices at optimal Hamming distance thresholds show TP, TN, FP, and FN counts, indicating strong detection capabilities even under complex paraphrasing attacks. Note that PI in the confusion matrices stands for Prompt Injection and is used to distinguish between attack and benign prompts.
(a) Precision-recall curves across attack variants.
(b) Confusion matrices at optimal thresholds.
Figure 6: SimHash baseline performance analysis across attack variants. Confusion matrices reveal growing FP and FN, particularly for paraphrasing scenarios where token-based similarity detection struggles.
3.2 BinaryShield Comparison with Baseline
To provide a rigorous performance analysis, we systematically compare BinaryShield against SimHash, a widely used locality-sensitive semantic hashing method (Manku et al., 2007). Both algorithms generate binary fingerprints and use Hamming distance for similarity measurement, making them directly comparable. This baseline is chosen to evaluate not only detection performance but also the privacy-preserving capabilities essential for cross-boundary threat intelligence sharing as SimHash also generates non-reversible binary fingerprints.
For this evaluation, we configure BinaryShield with a differential privacy parameter of $\alpha=2.0$, which provides meaningful privacy protection while maintaining strong detection performance. We note that BinaryShield with ModernBert generates 768-dimensional binary vector while SimHash produces 64-dimensional vector.
Our comparative analysis encompasses six distinct attack variants: minimal single-word modifications (V-1), progressive word changes (V-5, V-20) and extensive modifications in paraphrasing scenarios (V-Paraphrase) that fundamentally restructure prompts while preserving malicious intent. This progression allows us to assess how each method’s performance degrades as attack sophistication increases during threat correlation. Figure 5 and Figure 6 present comprehensive performance analysis through precision-recall curves (subfigures -a) and confusion matrices at optimal operating points (subfigures -b).
Figure 7: F1-score comparison between BinaryShield and SimHash across attack variants. BinaryShield consistently outperforms SimHash, particularly in complex paraphrasing scenarios, demonstrating superior robustness to semantic modifications.
For minimal attack modifications (V-1 to V-10), both algorithms demonstrate exceptional detection capabilities with near-perfect performance (Figure 7). At optimal thresholds, both methods achieve optimal F1-scores (Figure 7), establishing a strong baseline for more complex evaluations. For instance, in the V-10 scenario, SimHash produces 8 false positives and 40 false negatives at Hamming distance threshold ($\tau$) 19 (Figure 6), while BinaryShield generates only 4 false positives at threshold 276 (Figure 5).
As attack complexity increases to extensive word modifications (V-20), performance differences begin to emerge. BinaryShield maintains robust performance with an F1-score of 0.99 (precision: 0.99, recall: 0.98), while SimHash experiences measurable degradation to an F1-score of 0.96 (precision: 0.98, recall: 0.93). The 3.3 percentage point gap starts to illustrate BinaryShield’s superior stability as lexical modifications increase.
The most challenging scenario involves comprehensive paraphrasing attacks (V-Paraphrase), where prompts undergo fundamental restructuring while preserving malicious intent. Here, the performance differential becomes most pronounced: BinaryShield achieves an F1-score of 0.94 (precision: 0.92, recall: 0.95), while SimHash degrades significantly to an F1-score of 0.77 (precision: 0.75, recall: 0.80). This 17 percentage point gap in F1 score demonstrates BinaryShield’s fundamental advantage in threat intelligence that transcends surface-level linguistic variations.
Additionally, examining confusion matrices at optimal thresholds, provides additional insights into error patterns in Figures  5(b) and  6(b). BinaryShield demonstrates consistent performance as compared to SimHash. SimHash shows increasing error variability as attack complexity increases, struggling particularly to distinguish between semantically similar malicious content and unrelated benign prompts in paraphrasing scenarios.
Our empirical results demonstrate that BinaryShield provides superior robustness for prompt injection threat correlation compared to SimHash, with advantages becoming most pronounced for sophisticated attacks that represent complicated adversarial scenarios. The 17 percentage point F1-score gap in paraphrasing attacks, the most challenging and operationally relevant scenario, validates the effectiveness of BinaryShield over semantic hashing technique like SimHash. These performance improvements are achieved while simultaneously providing privacy guarantees essential for cross-boundary threat intelligence sharing. The superior performance on complex attacks is particularly significant, as sophisticated paraphrasing and extensive modification represent the most realistic attack patterns likely to be encountered in production environments where adversaries actively attempt to evade detection systems.
Summary. BinaryShield consistently outperforms SimHash in prompt injection threat correlation, especially under complex paraphrasing attacks, achieving an F1-score of 0.94 compared to SimHash’s 0.77. These results highlight BinaryShield’s superior robustness with enabling privacy, making it highly effective for cross-boundary threat intelligence sharing.
Figure 8: Privacy-utility trade-off for paraphrased prompt injection detection. (Left) BinaryShield F1 rises smoothly with privacy budget $\alpha$, dominating SimHash (privacy-preserving baseline) across most of the spectrum and approaching non-private dense embeddings. (Center) Precision lags recall at high noise, then rapidly converges after $\alpha\!\approx\!1$ as false positives collapse. Horizontal dashed lines: SimHash precision/recall. (Right) Confusion matrices at extreme settings: severe noise ($\alpha=0.25$) induces many false positives; higher $\alpha$ (33) sharply improves precision.
3.2.1 Empirical Analysis of Randomized Response Impact
We quantify the impact of randomized response mechanism (Warner, 1965) noise on detection utility by sweeping the per-bit privacy budget $\alpha\!\in\![0.25,3.0]$ (Algorithm 1, Step 4) on the most challenging paraphrase attack variant. For each $\alpha$, a bit in the 768-dimensional binary embedding is kept with probability $p=\tfrac{e^{\alpha}}{e^{\alpha}+1}$ and flipped otherwise, yielding an expected self-Hamming distortion (1p)d(1-p)d, where $d$ is the fingerprint dimension (here, $d=768$). At $\alpha=0.25$, $p=0.562$ and the mechanism flips $\approx 0.438\times 768\approx 336$ bits (i.e., the whole binary vector is approximately completely randomized). As shown in the top-right confusion matrix of Figure 8, at this extreme privacy setting there are 821 false positives and 97 false negatives, resulting in an F1-score of only 0.65. Such high noise levels provide maximal privacy but render the binary fingerprints nearly random (Figure 9), offering little practical utility for threat detection.
[…]
Randomized Response and Hamming Distance. To empirically validate BinaryShield’s privacy model, we conducted a controlled experiment measuring the actual Hamming distance introduced by the BinaryShield’s randomized response mechanism.
To this end, for 500 prompts from our evaluation dataset, BinaryShield generated 768-dimensional binary fingerprints and applied its differential privacy transformation by varying $\alpha$ values from 0.2 to 3.4. For each configuration, we computed the Hamming distance between original and noise-perturbed fingerprints, comparing observed distances against the theoretical expectation (1p)d(1-p)d. As a baseline reference, we generated 1,000 pairs of independent random binary vectors to establish the average distance for completely uncorrelated binary vectors. The mean Hamming distance for these independent vectors is $384.44\pm 13.65$, which serves as a baseline for comparison. Figure 9 demonstrates that our theoretical model precisely predicts the observed self-Hamming distortion across all privacy levels confirming BinaryShield’s alignment with theory.
Figure 9: Calibration of randomized-response noise. Empirical mean Hamming distances (blue points, $n=500$ each) between original and privatized 768-bit fingerprints tightly follow the theoretical curve (1p)d(1-p)d (solid black) across privacy budgets $\alpha$. The dashed red line shows the mean distance between independent random binary vectors (384), illustrating how increasing $\alpha$ moves BinaryShield smoothly from near-randomization to low-distortion states.
At $\alpha=0.2$, BinaryShield’s perturbed binary fingerprint reaches a Hamming distance of $346\pm 13.41$, nearly matching the random baseline. This results in lower F1 scores at small $\alpha$ (Figure 8).
As $\alpha$ increases, the Hamming distance decreases to $206.45\pm 13.08$ at $\alpha=1.0$. This decay directly correlates with the smooth utility recovery in Figure 8. The tight correspondence between theoretical predictions and observed measurements validates that BinaryShield’s privacy mechanism operates exactly as designed, providing operators with precise control over the privacy-utility trade-off through a single, well-characterized parameter $\alpha$. Another noticeable observation is that even at higher $\alpha$ values, as shown in Figure 9, the Hamming distance remains nonzero. This indicates that BinaryShield continues to provide privacy protection, since the nonzero Hamming distance ensures privacy is maintained, even as its utility approaches that of dense embeddings (Figure 8 and 9). In contrast, dense embeddings offer no privacy at all, so even a nonzero Hamming distance (i.e., some privacy) is fundamentally better than nothing, ensuring that sensitive information is not directly exposed.
Summary. BinaryShield exhibits the fundamental differential privacy trade-off: low $\alpha$ (strong privacy) yields near-random fingerprints, whereas increasing $\alpha$ (meaningful privacy) produces a smooth, monotonic F1 rise that quickly approaches non-private baseline performance.
Figure 10: Scalability of privacy-preserving threat correlation. Accuracy@k for BinaryShield vs non-private dense embeddings baseline across hybrid corpora (10K-100K). Accuracy remains flat with scale; BinaryShield retains 93% of baseline Accuracy@1 while providing privacy.
3.3 Real-World Deployment Analysis
While earlier sections validate BinaryShield supporior performance as compared to privacy preserving baseline (SimHash), we now turn to evaluating BinaryShield’s effectiveness in realistic cross-service threat intelligence scenarios. Specifically, it must be shown that privacy-preserving correlation remains effective when (i) malicious prompts are sparse relative to benign traffic and (ii) corpus size scales by an order of magnitude (iii) Privacy-utility trade-off is maintained, i.e., the accuracy gap between BinaryShield and non-private dense embeddings remains bounded, and (iv) efficiency is maintained in terms of computational and storage overhead.
We study this setting using hybrid corpora that interleave a (fixed) malicious prompt injection (V-Paraphrase) with large volumes of real user interactions from WildChat (Zhao et al., 2024), thereby emulating an enterprise log in which only a small fraction of entries are prompt injection attacks.
3.3.1 Large-Scale Threat Intelligence Scalability
The goal is to quantify how BinaryShield’s privacy-preserving fingerprint correlation accuracy behaves as the total corpus grows from 10K to 100K entries, and characterize the accuracy gap relative to a non-private dense embedding baseline. In this experiment, BinaryShield uses OpenAI’s text-embedding-3-large to generate dense embeddings.
Figure 11: Accuracy@k gap (Dense Embeddings - BinaryShield) across corpus sizes. Rapid narrowing for higher k demonstrate that binary quantization plus local DP introduce a constant, not compounding, utility cost.
For each target corpus size $C\in\{10\text{K},20\text{K},\dots,100\text{K}\}$, we construct a hybrid dataset by injecting a fixed malicious prompt set into CMC-M benign WildChat prompts (where $M$ is the number of distinct attack prompts). Each malicious prompt is issued as a query; success is recorded if its exact counterpart (or semantically identical variant, depending on $k$) appears within the top-$k$ candidates when ranked by Hamming distance over BinaryShield’s fingerprints, or cosine similarity for the dense baseline. We fix the local differential privacy parameter at $\alpha=2$ (different $\alpha$ values are also evaluated in Section 3.3.2).
[…]
Furthermore, Figure 11 shows that the Accuracy@k gaps between BinaryShield and the dense baseline with X-axis showing corpus sizes (10K to 100K) and Y-axis showing the gap at different top-k accuracy. The gap remains tightly bounded: it ranges from 5.31 to 6.15 percentage points (mean 5.66% points) at $k=1$, indicating that BinaryShield retains over 93% of the dense baseline’s accuracy even under tight local differential privacy constraints. Crucially, the gap does not significantly widen with corpus size. Furthermore, increasing $k$ rapidly closes the gap, as shown in Figure 10 and Figure 11. For instance at 100K entries, BinaryShield achieves 79.2% Accuracy@1, 90.8% Accuracy@3, and 92.9% Accuracy@5. On the same 100K corpus, the dense embeddings achieve 85.3% Accuracy@1, 94.7% Accuracy@3, and 96.5% Accuracy@5. The gap at $k=5$ is only 3.6% points, indicating that BinaryShield’s with privacy feature reaches close to the non-private baseline of dense embeddings. This confirms that the semantic signal preserved after (i) sign-only quantization and (ii) per-bit randomized response still remains sufficiently rich for practical threat correlation. For every corpus size, BinaryShield’s incremental gains from $k=1$ to $k=2$ average roughly +8 percentage points (Figure 10). This reflects that most residual misses at $k=1$ are near-boundary semantic neighbors retrievable with minimal expansion. Subsequent marginal gains ($k>2$) diminish smoothly. Figure 11 shows the gap curves monotonically compress as $k$ increases. The gap is attributable to two irreversible transformations: (a) many-to-one sign mapping (loss of magnitude information) and (b) randomized response bit flips calibrated by $\alpha=2$. That the gap does not drift upward with scale empirically supports that these transformations in BinaryShield behave as scale-neutral perturbations rather than compounding sources of semantic erosion. In an operational setting where an analyst (or automated response engine) can examine the top 3–5 correlated candidates, BinaryShield delivers privacy-preserving threat intelligence with accuracy comparable to a fully non-private dense embedding pipeline. This constitutes a strong utility guarantee under stringent local privacy constraints and validates BinaryShield’s suitability for multi-service deployment where corpus sizes routinely exceed tens of thousands of daily log entries.
Summary. BinaryShield achieves over 93% accuracy of dense embeddings (non-private baseline) and remains scalable in practical deployment settings with randomized response mechanism, demonstrating robust privacy-preserving threat correlation at enterprise scale.
3.3.2 BinaryShield’s Privacy Parameter Performance at Scale
Figure 12: Privacy-utility trade-off for BinaryShield across $\alpha$ values on 50K hybrid corpus. Accuracy@1 (blue line) and Accuracy@5 (dotted black line) show phase transition behavior as privacy budget increases. The non-private dense embedding baseline is shown as horizontal green and orange lines at 86.04% (Accuracy@1) and 96.56% (Accuracy@5) respectively.
The relationship between privacy protection and threat detection utility at scale represents a fundamental tension in the cross-boundary threat intelligence system. Similar to Section 3.2.1, to characterize this trade-off precisely, we evaluate BinaryShield’s performance across a spectrum of privacy budgets ($\alpha$) on a fixed 50K-entry hybrid corpus containing both malicious prompts and benign WildChat interactions (Section 3.3.1). The privacy parameter $\alpha$ directly controls the randomized response mechanism through the bit preservation probability $p=e^{\alpha}/(e^{\alpha}+1)$, with the corresponding bit flip probability being 1p1-p. Smaller $\alpha$ values provide stronger privacy guarantees by increasing the probability of bit flips in the binary fingerprint, while larger values preserve more semantic information at the cost of reduced privacy protection.
Our experimental results, visualized in Figure 12, reveal a sharp phase transition in detection accuracy as the privacy budget increases from highly restrictive to meaningful protective settings. The X-axis represents the privacy parameter $\alpha$, while the Y-axis shows correlation accuracy (Accuracy@1 and Accuracy@5) for both BinaryShield and the non-private dense embedding baseline.
At extreme privacy levels ($\alpha=0.1$), the bit flip probability reaches 0.475, meaning nearly half of all bits are randomly inverted, effectively reducing fingerprints to near-random noise. Consequently, Accuracy@1 drops to 0%, rendering threat correlation impossible. This represents the theoretical limit where maximal privacy completely eliminates utility. As we relax the privacy constraint the bit flip probability decreases. With $\alpha=0.75$ (bit flip probability 0.32), Accuracy@1 rises to 14.17% and Accuracy@5 to 26.15%, marking the onset of effective threat detection. At $\alpha=1.0$ BinaryShield to reach 44.06% Accuracy@1 and 64.90% Accuracy@5. While this is roughly half the non-private baseline (86.04% and 96.56%), it shows that meaningful threat correlation is feasible even under strong privacy constraints. Between $\alpha=1.0$ and $\alpha=2.0$, accuracy improves smoothly with meaningful privacy and utility tradefoff. At $\alpha=1.25$, Accuracy@1 reaches 62.40% and Accuracy@5 82.81%. At $\alpha=1.5$, Accuracy@1 is 72.60% and Accuracy@5 is 89.90%, retaining 84.4% of baseline Accuracy@1.
The results show a smooth transition in the privacy-utility trade-off, especially between $\alpha=0.75$ and $\alpha=2.0$. For significantly high-privacy needs, $\alpha=1.5$ gives strong privacy. Since organizations have already strict security policies, for internal sharing, relaxing privacy to $\alpha=2.0$$2.5$ yields near-optimal utility with realistic meaningful privacy protection. Overall, BinaryShield potentially can operate in a wide range of privacy-utility trade-offs.
Summary. BinaryShield empowers organizations to balance privacy and utility as needed, consistently delivering strong accuracy and realistic privacy protection. The organizations can flexibly tune the privacy-utility trade-off based on their operational needs, regulatory requirements, and threat detection scenarios.
3.3.3 BinaryShield Computational and Storage Efficiency Analysis
The operational viability of any threat intelligence system fundamentally depends on its ability to process vast quantities of data with minimal latency and storage overhead. Historical lessons from traditional cybersecurity infrastructure underscore this requirement. One of the major reason that signature-based ecosystems (antivirus, IDS) scaled is because compact hashes (e.g., MD5, fuzzy-hash families). Similarly, modern LLM services face an even more daunting challenge, processing billions of queries daily across distributed infrastructure. For BinaryShield to enable practical cross-boundary threat intelligence, it must achieve orders-of-magnitude improvements in both computational speed and storage efficiency compared to non-private dense embeddings. Every extra byte per fingerprint multiplies into material storage cost, and every millisecond of per-fingerprint scan latency stretches incident response time.
Figure 13(a) reveals the stark computational advantages of BinaryShield’s binary fingerprints over dense embeddings. At a corpus size of 10K entries, dense embeddings require 0.87 seconds for similarity search of 968 prompt injections, while BinaryShield completes the same operation in just 0.032 seconds, a 26.7x speedup. This performance gap widens dramatically as corpus size increases. At 100K entries, dense embeddings demand 14.52 seconds, rendering real-time threat correlation infeasible for high-volume services. In contrast, BinaryShield maintains sub-second performance at 0.38 seconds, achieving a 38.1x speedup. The computational efficiency stems from the fundamental difference in similarity computation. Dense embeddings require floating-point dot products or cosine similarity calculations across thousands of dimensions, operations that scale poorly even with optimized linear algebra libraries. BinaryShield’s binary fingerprints enable Hamming distance computation through simple XOR operations.
(a) Search time comparison between non-private dense embeddings and BinaryShield across corpus sizes. Note the logarithmic scale on the Y-axis. Dense embeddings exhibit super-linear growth reaching 16.3 seconds at 80K entries, while BinaryShield maintains sub-second performance throughout, demonstrating its suitability for real-time threat correlation at enterprise scale.
(b) Storage size comparison demonstrating BinaryShield’s 64x reduction compared to dense embeddings. Dense embeddings scale from 256.88 MB at 10K entries to 2,366.25 MB at 100K entries, while BinaryShield requires only 4.01 MB to 36.97 MB respectively, enabling cost-effective long-term threat intelligence retention.
Figure 13: Computational and storage efficiency analysis of BinaryShield versus dense embeddings.
The storage requirements present an equally compelling case for BinaryShield’s architecture, as illustrated in Figure 13(b). Dense embeddings consume 256.88 MB for 10K entries, growing linearly to 2,366.25 MB at 100K entries. This represents a substantial infrastructure burden when organizations must maintain threat intelligence databases spanning billions of historical queries across multiple compliance boundaries. BinaryShield reduces storage requirements by a factor of 64, requiring only 4.01 MB for 10K entries and 36.97 MB for 100K entries. This 64x reduction directly translates from the binary quantization process, where each embedding dimension is compressed from a floating-point value to a single bit, followed by the addition of noise through randomized response that preserves the binary nature.
The practical implications of these efficiency gains extend beyond raw performance metrics (Section  3.2 and  3.3.1). Consider an organization processing 100 million queries daily across ten compliance boundaries. With dense embeddings, maintaining a rolling 30-day threat intelligence window would require approximately 7.1 TB of storage per boundary and dedicated GPU clusters for similarity search making it near to impractical choice. BinaryShield reduces this to 111 GB per boundary, fitting comfortably in memory on commodity servers while enabling CPU-based similarity search at wire speed. This efficiency democratizes threat intelligence capabilities, allowing even resource-constrained services to participate in collaborative defense without significant infrastructure investment.
Furthermore, the computational efficiency enables new operational capabilities previously infeasible with dense embeddings. Security teams can now perform retrospective threat hunting across months of historical data in minutes rather than hours. The reduced storage footprint also facilitates comprehensive threat intelligence archival, enabling long-term trend analysis and attribution of persistent attack campaigns that evolve over extended periods.
Summary. BinaryShield’s computational and storage efficiency fundamentally transforms cross-boundary threat intelligence economics. By reducing search latency up to 38x and storage requirements up to 64x while maintaining 93% of non-private baseline top-1 accuracy with privacy, BinaryShield enables enterprise-scale threat correlation. This positions BinaryShield as a practical foundation for collaborative threat intelligence systems against prompt injection attacks that respect regulatory boundaries while enabling effective cross-service intelligence sharing.
4 Related Work and Discussion
Prompt injection attacks pose significant risks to LLMs by embedding covert instructions within benign-looking input data to manipulate model outputs. A number of recent studies have explored both the attack methods and potential defenses against these vulnerabilities.
4.1 Attacks
[…]
4.2 Attack Defenses
Recent work proposes a variety of defense strategies against prompt injection attacks. These approaches differ in their methods and operating settings, ranging from inference-time countermeasures to white-box techniques and architectural safeguards.
Yuan et al. (2024) propose RigorLLM framework that enhances content moderation by (a) enriching training data with diverse adversarial examples using constrained energy-based generation, (b) appending an optimized “safe suffix” to incoming prompts to mitigate adversarial modifications, and (c) fusing predictions from a robust KNN classifier and a fine-tuned LLM to reliably detect harmful content.  Hines et al. (2024) build on the work of  Yi et al. (2025) [^1] We cite the peer-reviewed conference version. An earlier preprint is available on arXiv. to address indirect prompt injection attacks by proposing a defense strategy based on spotlighting. This approach comprises a set of prompt-engineering techniques that transparently indicate the provenance of each input segment to assist LLM while generating response. The methods operate exclusively at inference time and do not require access to the model’s internals, making them readily applicable to API-based models.  Debenedetti et al. (2025) propose CaMeL, a defense mechanism that explicitly detects and prevents prompt injection attacks by designing a secure execution layer around LLM operations. The work sits at the intersection of generative AI and software security, bringing established principles from the latter to protect emerging systems. Abdelnabi et al. (2025a) proposes TaskTracker, white-box defense, that centers on tracking the model’s internal activations before and after processing external text. By calculating the change (or “delta”) between these activations, the system can flag when the model’s task representation has been altered by injected instructions. This drift is used as a proxy to signal potential prompt injections.  Liu et al. (2025) introduces DataSentinel that detects prompt injection attacks by fine-tuning a language model with a game-theoretic mini-max optimization. It uses a detection instruction embedding a secret key to verify the integrity of incoming data. Beurer-Kellner et al. (2025) propose a six architectural design patterns (action-selector, plan-then-execute, map-reduce, dual LLM, code-then-execute, and context-minimization) that secure LLM agents against prompt injection attacks. These patterns restrict an agent’s ability to process untrusted input with some trade-offs in agent flexibility.
Overall, most existing defenses are probabilistic and can be bypassed (Debenedetti et al., 2025; Costa et al., 2025). Retrospective systems are needed to identify attackers, yet cross-service threat intelligence for large-scale prompt injection attacks correlation remains unexplored.
4.3 Discussion
Our study has limitations similar to those in any research. Synthetic paraphrasing and word substitution may yield varying results with different LLMs and prompts. The effectiveness of the BinaryShield’s PII redaction module depends on the accuracy of detection methods. Residual PII may remain in prompts. Our results rely on specific embedding models, ModernBert (open-source) (Warner et al., 2025; Gill et al., 2025) and OpenAI text-embedding-3-large (proprietary). Thus, performance may vary with other models. We determined optimal Hamming distance thresholds through search on our datasets. However, these thresholds may not generalize to other attack distributions or organizational contexts. All experiments contain English prompts. Future work should explore cross-lingual attack detection and domain-specific technical prompts. We depend on computational non invertibility and differential privacy guarantees. The evaluation does not assess resistance to reconstruction attacks, which represent an orthogonal concern to current focus of the paper.
Noise can be added to the fingerprint through various differential privacy mechanisms (Warner, 1965; Wang et al., 2017; Bhowmick et al., 2018; Erlingsson et al., 2014; Feyisetan and Kasiviswanathan, 2021; Vadhan and Wang, 2021; Carvalho et al., 2021; Bollegala et al., 2023; Xie et al., 2024; Meehan et al., 2022; Asi et al., 2022; Du et al., 2023; Bittau et al., 2017; Fanti et al., 2016; Hsu et al., 2014; Near and Abuah, 2021; Fioretto et al., 2024; Dwork et al., 2014; Balle and Wang, 2018; Ding et al., 2017; Hu et al., 2024). BinaryShield utilizes foundational work of Randomized Response mechanism (Warner, 1965). Assessing the impact of other DP mechanisms and global differential privacy on the overall privacy guarantees is an important direction for future research. We did not formulate the privacy utility trade off as an optimization problem. Readers may consult existing literature on this topic (Near and Abuah, 2021; Wang et al., 2017; Dwork et al., 2014; Fioretto et al., 2024). However, in our experiments, we empirically evaluated conditions ranging from extreme privacy, where accuracy approaches zero, to moderate and weak privacy. Each organization has specific privacy policies, and we do not recommend a particular differential privacy budget. Organizations should work with their privacy teams to select an appropriate budget based on their guidelines. Note we do not claim any novelty in the individual techniques used in BinaryShield. Our primary contribution is the proposal of the first threat intelligence system for prompt injection threat correlation. We integrate insights from data anonymization, natural language processing, quantization, and differential privacy to design a practical and deployable threat intelligence system. This work constitutes an initial step in the field, and we hope it will inspire further research in cross-service threat intelligence for LLM misuse.
5 Conclusion
The proliferation of LLMs across enterprise services has created an unprecedented security challenge: organizations operate multiple LLM services with billions of daily queries, yet regulatory boundaries prevent these services from sharing critical threat intelligence about prompt injection attacks. This fragmentation leaves each service vulnerable to attacks that may have already been detected elsewhere within the same organization, creating a fundamental gap in collective defense capabilities.
BinaryShield addresses this critical vulnerability by introducing the first privacy-preserving fingerprinting mechanism specifically designed for prompt injection threat intelligence. Through a carefully orchestrated pipeline combining PII redaction, semantic embedding, binary quantization, and randomized response, BinaryShield generates non-invertible fingerprints that can be safely shared across compliance boundaries while preserving the semantic characteristics necessary for effective threat correlation. When compared to SimHash, the privacy-preserving baseline, BinaryShield demonstrates superior robustness. The operational efficiency gains are equally compelling.
The implications of this work extend beyond immediate operational benefits. As LLM-based systems become critical infrastructure across finance, healthcare, and government services, the attack surface for prompt injection continues to expand exponentially. The emergence of autonomous agents and MCP servers further amplifies these risks, potentially enabling attacks that cascade from text manipulation to arbitrary code execution and system compromise. BinaryShield establishes the foundational infrastructure for collaborative defense against these evolving threats, creating a pathway toward industry-wide threat intelligence feeds analogous to existing malware signature ecosystems.